A scuba diver dives with a tank containing 5L of air. Twenty-two percent of the air is oxygen. If the diver produces 1.62 mol carbon dioxide for every 2.05 mol of oxygen he/she inhales, what is the volume (at STP) of carbon dioxide produced if 4.2L of air are consumed at 202.6 KPa and 14.7oC
1
Expert's answer
2019-10-21T07:16:52-0400

n(O2)=4.2*0.22*202.6/(8.314*(14.7+273.15))=0.078 mol

n(CO2)=0.078*1.62/2.05=0.0618 mol

V(CO2)=0.0618*8.314*(14.7+273.15)/202.6=0.73 l

Answer: 0.73 l.